# TALC: Time-Aligned Captions for Multi-Scene Text-to-Video Generation

### v3

@@ 8 Acknowledgement
Appendix A Limitations
While there are numerous text-to-video generative models, our work focuses on diffusion-based T2V models due to their higher visual quality and better latency in comparison to autoregressive models [19]. In our work, we argue that the text conditioning mechanism is unified across diffusion T2V models, and thus TALC framework should generalize to various model architectures. Specifically, we perform our experiments with ModelScope, a latent diffusion model, and Lumiere, a diffusion model that operates in the raw video space. In the future, it will helpful to adopt TALC for multi-scene video generation in novel training and model design paradigms.
In this work, we evaluate the multi-scene video generation capabilities uptill four scenes. This is at par with the number of scenes in prior multi-scene video generation works [18, 23]. In our experiments, we further show that the performance of the models degrade as we generate more scenes, and hence, we limited ourselves to four scenes. A relevant future direction would be to expand the quality of multi-scene video-text data beyond four scenes and finetune T2V models using TALC framework.
In addition, we perform human annotations using Amazon Mechanical Turkers (AMT) where most of the workers primarily belong to the US and Canada (hourly rate - \$$\$18$ per hour). Hence, the human annotations in this work do not represent the diverse demographics around the globe. As a result, our human annotations reflect the perceptual biases of the annotators from Western cultures. In the future, it will be pertinent to assess the impact of diverse groups on our human evaluations. Finally, we acknowledge that the text-to-video generative models can perpetrate societal biases in their generated content [45, 46]. It is critical that future work quantifies this bias in the generated videos and provides methods for safe deployment of our models.
Appendix B Detailed Related Work
Text-to-Video Generative Modeling.
The field of text-to-video (T2V) synthesis has significantly evolved from its inception with models like VGAN [2] and MoCoGAN [41], leveraging the foundational technologies of GANs [42] and VAEs [43] to produce concise, single-scene videos. The narrative depth was further expanded through transformer-based architectures such as CogVideo [47] and VideoGPT [48], enhancing the complexity of video content yet remaining within the confines of single scenes. The advent of diffusion models, exemplified by Imagen Video [44], marked a notable advancement in T2V synthesis. Despite these strides, the challenge of creating multi-scene videos that reflect the complexity of the physical world [1, 2, 3] remains. Our work, TALC, extends the capabilities of T2V models to multi-scene storytelling, filling a crucial gap in the synthesis landscape.
Image-to-Video Animation.
The exploration of multi-scene video generation, innovative methods such as Lumiere [11] and Make-a-Video [49] have employed a two-step process, transforming text to images and then animating these images into videos. While these approaches have advanced visual quality, they often fall short in weaving seamless multi-scene narratives. This limitation is echoed in the work of Emu Video [7], which underscores the difficulty of achieving narrative coherence across multiple scenes. TALC focuses on direct generation of multi-scene narratives from textual prompts aiming for a narrative flow and visual consistency across scenes.
Multi-Scene Video Generation.
The pursuit of multi-scene T2V synthesis has been furthered by recent innovations like Phenaki [18] and Stable Video Diffusion [9], which have explored new frontiers in video generation from textual prompts and the scaling of latent diffusion models, respectively. Additionally, Dreamix [50] and Pix2Video [51] have broadened the scope of diffusion models, applying them to video editing and animation. Despite these advancements, the task of generating videos that convey coherent narratives across multiple scenes remains formidable, highlighted by recent works such as VideoPoet [19], ModelScope [10] and Make-A-Scene [52]. TALC tackles this task and offers a framework produces videos spanning multiple scenes. We also introduce nuanced evaluation approach. This approach integrates both automated assessments and human evaluations to rigorously gauge the quality and narrative coherence of the generated content, evaluating text adherence, object consistency and background consistency, contributing to the ongoing refinement of T2V synthesis.
Appendix C Qualitative comparison between baselines and TALC
We provide an qualitative example to showcase the benefit of TALC over merging captions and merging videos baselines in Figure 6.
Figure 6: Examples of multi-scene video generation baselines. (a) Generating video on the merged descriptions, leads to a poor text-video alignment. (b) Generating videos for the individual text descriptions and concatenate them temporally, leads to a lack of visual consistency. (c) Our approach (TALC) enhances the scene-level text-video alignment and maintains visual consistency.
Appendix D Prompt for Multi-Scene Caption Generation
@@
Appendix E Source datasets for multi-scene video-text data creation
@@ I.1 Task Prompts
Merging captions (Base) 95.9 43.8 69.9 99.5 21.5 60.5 81.9 32.0 56.9
Merging videos (Base) 69.4 71.5 70.5 71.2 82.3 76.7 57.9 58.7 58.3
TALC (Base) 93.6 48.3 71.0 99.3 57.3 78.3 81.4 36.1 58.8
Merging captions (F.T.) 94.2 57.1 75.7 94.9 31.2 63.1 50.8 24.3 37.5
TALC (F.T.) 93.3 62.6 77.9 98.9 76.3 87.6 79.7 47.9 63.8
@@ I.2 Number of Scenes
Merging captions (Base) 93.2 34.4 63.8 92.5 33.0 62.7 87.4 29.9 58.6
Merging videos (Base) 66.7 69.9 68.3 65.2 71.9 68.6 63.5 70.7 67.1
TALC (Base) 92.6 54.4 73.5 89.8 48.0 68.9 87.3 39.3 63.3
Merging captions (F.T.) 87.7 45.7 66.7 83.2 39.5 61.4 60.0 27.4 43.7
TALC (F.T.) 88.5 66.6 77.5 90.7 64.1 77.4 87.8 56.1 71.9
@@
Appendix M Visual quality of the generated videos.
Method Quality
Merging Captions (Base) 80.5
Merging Videos (Base) 86.5
TALC (Base) 84.5
Merging Captions (F.T.) 63.4
TALC (F.T.) 83.3
We compare the visual quality of the generated videos using human evaluation in Table 7. We find that the visual quality of videos generated from the base model ranges from $80.5-86.5$ using the baselines and TALC framework. However, we observe that the visual quality of generated videos is quite poor for the model finetuned with merging captions with a score of $63.4$ points. This highlights that finetuning a T2V model with multi-scene video-text data by naively merging the scene-specific descriptions in the raw text space leads to undesirable artifacts in the generated video. Finally, we find that the TALC-finetuned model ($83.3$) achieves a video quality score similar to that of the TALC-base model ($84.5$), indicating that our finetuning data preserves the visual quality observed during the model’s pretraining.
Appendix N Single-Scene Video Generation
@@
Figure 12: Examples of videos generated by Time-Aligned Captions (TALC)

### v4

@@ 7 Acknowledgement
Appendix A Related Work
Text-to-Video Generative Modeling:
Diffusion models like Imagen Video [21] represent a significant advancement in T2V synthesis, yet generating multi-scene videos that realistically capture the complexity of the physical world [1, 47, 10] remains challenging. Recent research has attempted longer video generation, but limitations persist. Phenaki [45] targets arbitrary-length videos with a focus on temporal coherence, though its evaluation is constrained by the lack of publicly available code. VideoDirectorGPT [30], DirecT2V [23], and Free-bloom [24] employ zero-shot approaches for multi-scene generation, contrasting with our fine-tuned method for enhanced performance. Gen-L-Video [48] uses iterative denoising to create consistent videos by aggregating overlapping short clips, but it typically generates videos with a maximum of two scenes, whereas our method supports up to four distinct scenes. StreamingT2V [19] employs an auto-regressive, streaming-based approach for continuous video generation. While these methods contribute to long video generation, they often struggle with maintaining visual consistency and adherence to multi-scene textual descriptions, which are critical for storytelling.
Multi-Scene Video Generation:
Efforts like Phenaki [45] and Stable Video Diffusion [7] push the boundaries of text-driven generation by scaling latent diffusion models. Dreamix [34] and Pix2Video [12] utilize diffusion models for video editing and animation, while methods like MCVD [46] and VDT [32] focus on improving temporal consistency in longer videos. Despite these advances, generating multi-scene videos that accurately reflect complex narratives with high visual fidelity remains difficult, as shown by ongoing research in projects like VideoPoet [28], ModelScope [49], and Make-A-Scene [17]. TALC addresses the challenges of multi-scene video generation by enhancing visual consistency and text adherence across scenes. Unlike zero-shot methods, TALC leverages fine-tuning on a curated multi-scene video dataset. We also propose a comprehensive evaluation protocol, combining human evaluation and automated assessments using GPT-4V, to ensure robust narrative coherence in the generated videos.
Appendix B Evaluator
Human Evaluation. Here, we use the annotators from Amazon Mechanical Turk (AMT) to provide their judgements for the generated videos. Specifically, we choose the annotators that pass a preliminary qualification exam. Subsequently, they assess the multi-scene generated videos along the dimensions of entity and background consistency, text adherence, and visual quality. For each metric, the multimodal model assigns one of three possible response $\{yes=1,partial=0.5,no=0\}$. For instance, $yes$ for the entity consistency metric implies that the video frames sampled from the generated video have consistent appearance of the entity described in the multi-scene script. We present the screenshot of the UI in Appendix §H.
Automatic Evaluation. Here, we utilize the capability of a large multimodal model, GPT-4-Vision [35], to reason over multiple image sequences. First, we sample four video frames, uniformly, from each scene in the generated video (e.g., 8 videos frames for two-scene video). Then, we prompt the multimodal model with the temporal sequence of video frames from different scenes and the multi-scene text description. The model is instructed to judge the generated videos for visual consistency and text adherence, similar to the human evaluation. In this work, we do not utilize any existing video-text alignment models [52, 5] for evaluating text adherence as they are trained on single-scene video-text datasets. We find that the agreement between the automatic evaluation and human evaluation is $77\%$. We present the automatic evaluation prompt in Appendix §G.
Appendix C Multi-scene video data sources
Appendix D Visual quality of the generated videos.
Method Quality
FreeNoise 80
Merge captions 80.5
Merge videos 86.5
Finetuning w/ Merge captions 63.4
TALC 84.5
Finetuning w/ TALC 83.3
We compare the visual quality of the generated videos using human evaluation in Table 4. We find that the visual quality of videos generated from the base model ranges from $80-86.5$ using the baselines and TALC framework. However, we observe that the visual quality of generated videos is quite poor for the model finetuned with merging captions with a score of $63.4$ points. This highlights that finetuning a T2V model with multi-scene video-text data by naively merging the scene-specific descriptions in the raw text space leads to undesirable artifacts in the generated video. Finally, we find that the TALC-finetuned model ($83.3$) achieves a video quality score similar to that of the TALC-base model ($84.5$), indicating that our finetuning data preserves the visual quality observed during the model’s pretraining.
Appendix E Prompt for Multi-Scene Caption Generation
@@ I.1 Task Prompts
Merge captions 95.9 43.8 69.9 99.5 21.5 60.5 81.9 32.0 56.9
Merge videos 69.4 71.5 70.5 71.2 82.3 76.7 57.9 58.7 58.3
F.T. w/ merge captions 94.2 57.1 75.7 94.9 31.2 63.1 50.8 24.3 37.5
TALC 93.6 48.3 71.0 99.3 57.3 78.3 81.4 36.1 58.8
F.T. w/ TALC 93.3 62.6 77.9 98.9 76.3 87.6 79.7 47.9 63.8
@@ I.2 Number of Scenes
Merge captions 93.2 34.4 63.8 92.5 33.0 62.7 87.4 29.9 58.6
Merge videos 66.7 69.9 68.3 65.2 71.9 68.6 63.5 70.7 67.1
F.T. w/ merge captions 87.7 45.7 66.7 83.2 39.5 61.4 60.0 27.4 43.7
TALC 92.6 54.4 73.5 89.8 48.0 68.9 87.3 39.3 63.3
F.T. w/ TALC 88.5 66.6 77.5 90.7 64.1 77.4 87.8 56.1 71.9
@@
Appendix M Single-Scene Video Generation
Figure 10: Examples of videos generated by TALC with Lumiere, FreeNoise and Gen-L-Video. The videos generated using TALC with Lumiere exhibited smooth motion, seamless transitions between text prompts, and high overall quality. In contrast, videos created with FreeNoise were of lower quality, with notably less motion. For example, in the second video, the background remained static except for darkening, and the jeep did not appear to be in motion. Videos generated with Gen-L-Video demonstrated less smooth transitions between prompts and, at times, showed inconsistencies in object representation, as seen in the first video where the man’s appearance changed repeatedly.
Figure 11: Qualitative examples of videos generated by TALC with ModelScope. Videos generated with TALC allow for the creation of videos where a single object performs multiple actions while maintaining visual consistency. Furthermore, TALC facilitates the generation of videos with diverse scenes, as demonstrated in the final example.
@@
Appendix N FreeNoise failure cases
In this work, we observe that the state-of-the-art method FreeNoise [38] does not perform well on our multi-scene prompts. We provide some qualitative examples to show its inability to introduce new content and limited motions as multiple scenes are requested in Figure 12, 13, and 14.
Figure 12: FreeNoise generation for the prompt: "A wolf howls at the moon in a dense forest.; The wolf prowls stealthily through the underbrush, eyes gleaming.; The wolf catches a scent and follows it eagerly.; The wolf emerges victorious, holding a fresh kill." Explanation: The generated scene shows a wolf standing in a forest, but the key actions described in the prompt are missing. The wolf does not prowl, its eyes are not gleaming, it does not follow any scent, and no kill is presented. This indicates a significant gap in the model’s ability to convey the sequential and dynamic nature of the narrative.
Figure 13: FreeNoise generation for the prompt: "A chimpanzee swings effortlessly through the forest canopy.; The chimpanzee gathers fruit, chattering with companions." Explanation: Instead of depicting the chimpanzee swinging through the canopy, the generated scene shows it sitting on a tree branch and then on the ground. Furthermore, the scene lacks the critical elements of the prompt, such as the presence of fruits and companions, which are essential for conveying the intended narrative.
Figure 14: FreeNoise generation for the prompt: "A tiger prowls through dense jungle undergrowth stealthily.; The tiger ambushes a deer, swift and deadly." Explanation: The scene fails to show the tiger ambushing a deer as described. Instead, the tiger is depicted standing calmly, and no deer is visible. Additionally, a second tiger appears unexpectedly from the first tiger’s body, abruptly changing directions, which introduces further inconsistencies and detracts from the intended narrative.